# Generic groups and the weak amalgamation property

### v2

@@ 2.3 Nilpotent group of prime exponent $p$ with $p>2$
Proof. We need some material concerning amalgamation bases in 2-step nilpotent groups, see Theorems 3.3 - 3.6 in [13]. It can be deduced from these theorems that a 2-step nilpotent group $G$ of exponent $p$ is an amalgamation base if and only if $G^{\prime}=C(G)$. Below we only need sufficiency of this and the latter can be shown easier. Indeed, consider $G\in\mathbb{G}_{2_{nil}p}$ with $G^{\prime}=C(G)$. Note that any embedding of $G$ into any $H\in\mathbb{G}_{2_{nil}p}$ preserves every unary predicate $P$ with the property $H^{\prime}\leq P\leq C(H)$ where $P$ on $G$ is defined by $P=G^{\prime}=C(G)$. By Lemmas 2.2 and 2.3 of [1] the class $\mathbb{G}_{2_{nil}p}$ in the language extended by $P$ with $H^{\prime}\leq P\leq C(H)$, has the amalgamation property.
It now suffices to verify that every finite group $G_{0}\in\mathbb{G}_{2_{nil}p}$ extends to a finite group $G_{1}\in\mathbb{G}_{2_{nil}p}$ with $G^{\prime}_{1}=C(G_{1})$. This can be again deduced from the amalgamation result of [1]. Indeed, consider $G_{0}$ with the predicate $P(G_{0})=C(G_{0})$. Since $P(G_{0})$ is a vector space over $\mathbb{F}_{p}$ we decompose $P(G_{0})=G^{\prime}_{0}\oplus\langle e_{1},\ldots,e_{k}\rangle$ where $e_{1},\ldots,e_{k}$ is a basis of the complement of $G^{\prime}_{0}$. Now take a finite $H\in\mathbb{G}_{2_{nil}p}$ such that $H^{\prime}=C(H)=\langle e_{1},\ldots,e_{k}\rangle$. The amalgamation process described in Lemmas 2.2 and 2.3 of [1] applied to $(G_{0},P)$ and $(H,P)$ gives the result. $\Box$
We do not know if the statement of this proposition holds for $c$-step nilpotent groups where $c>2$. The argument in the beginning of the proof together with [2] proves that if a group $G\in\mathbb{G}_{c_{nil}p}$ satisfies $\Gamma_{i}(G)=Z_{c+1-i}(G)$ for all $1<i\leq c$ then $G$ is an amalgamation base. Is it possible to embed every $H\in\mathbb{G}_{c_{nil}p}$ into $G$ with this property?
@@ 3 Strong undecidability implies non-WAP
In this section we will give several group-theoretic properties where the corresponding space of enumerated groups does not have generics. Throughout the section we are under the setup of one of the cases I - II of Section 1.2. We will assume below that $Th(M)$ is $\omega$-categorical and $T$ is a universal theory including group axioms for the sort $\mathsf{Gp}$.
@@ 3.1 Completeness and Kuznetsov’s theorem
Assume that $D(\bar{c})\in\mathcal{B}_{T}$ and an extension $D^{\prime}(\bar{c}^{\prime})\in\mathcal{B}_{T}$ is t-isolating for $\bar{c}$. Let $(M,{\bf\bar{r}})\in[D^{\prime}(\bar{c}^{\prime})]$. Assume that $D^{\prime}(\bar{c}^{\prime})$ is a computably enumerable set. Then in $(M,{\bf\bar{r}})$ the elements of $\bar{c}$ which belong to the group sort $\mathsf{Gp_{\bar{r}}}$ generate a recursively presented group with decidable word problem.
Proof. We may assume that $\bar{c}$ contains all elements of the sort $\mathsf{Gp_{\bar{r}}}$ which are distinguished in $(M,{\bf\bar{r}})$ as constants. Let us extend $D(\bar{c})$ by all equations and inequations with parameters from $\bar{c}$ which can be logically deduced from $D^{\prime}(\bar{c}^{\prime})$. The obtained set, say $\Sigma$, is computably enumerable. The set of group words $w$ with $(w=1)\in\Sigma$ forms a computable presentation of $\langle\bar{c}\cap\mathsf{Gp_{\bar{r}}}\rangle$. Remember that this presentation does not depend on the expansion $(M,{\bf\bar{r}})$.
The deciding algorithm is as follows. Let $w$ be a group word in the alphabet $\bar{c}\cap\mathsf{Gp_{\bar{r}}}$. We verify $(w=1)\in\Sigma$ using the corresponding enumeration. The algorithm stops when this procedure gives $w=1$ or $w\not=1$ in $\Sigma$. If the first case does not hold then in any member of $[D^{\prime}(\bar{c}^{\prime})]$ the group $\langle\bar{c}\cap\mathsf{Gp_{\bar{r}}}\rangle$ satisfies $w\not=1$. If $w\not=1$ does not belong to $\Sigma$ then $Th(M,\bar{c}^{\prime})\cup D^{\prime}(\bar{c}^{\prime})\cup T\cup(w=1)$ is consistent. This contradicts $\omega$-categoricity of $Th(M,\bar{c}^{\prime}))$. We see that the second case holds. $\Box$
The following theorem will be the main tool for proving absence of generics in $\mathcal{G}_{T}$.
@@
Assume that $D(\bar{c})\in\mathcal{B}_{T}$ (where $\bar{c}$ includes all distinguished elements of $M$) satisfies the property that every extension $D^{\prime}(\bar{c}^{\prime})$ is computably enumerable and for every $(M,{\bf\bar{r}})\in[D(\bar{c})]$ the group $\langle\bar{c}\cap\mathsf{Gp_{\bar{r}}}\rangle$ (defined in $(M,{\bf\bar{r}})$) has undecidable word problem.
@@
Proof. Assume that an extension $D^{\prime}(\bar{c}^{\prime})$ witnesses WAP. By Proposition 3.1, $D^{\prime}(\bar{c}^{\prime})$ is not t-isolating for $\bar{c}$. This means that there are $(M,{\bf\bar{r}})$, $(M,{\bf\bar{r}^{\prime}})\in[D^{\prime}(\bar{c}^{\prime})]$ and a word $w(\bar{c}_{0})$ with $\bar{c}_{0}\subseteq\bar{c}\cap\mathsf{Gp}$ such that $(M,{\bf\bar{r}})\models w(\bar{c}_{0})=1$ but $(M,{\bf\bar{r}^{\prime}})\models w(\bar{c}_{0})\not=1$. We see that the two extensions of $D^{\prime}(\bar{c}^{\prime})$ by $w(\bar{c}_{0})=1$ and by $w(\bar{c}_{0})\not=1$ (domains of these extensions are enriched by values of subwords appeared in $w$) cannot be amalgamated over $D(\bar{c})$. $\Box$
We now give several applications of Theorem 3.2 for concrete spaces of enumerated groups. In these applications we are under Scenario I of Section 1.2. In particular the structure $M$ is just $(\omega,1)$.
@@
In Proposition 5.2.13 of [6] the authors present a construction (of A. Darbinyan) of a finitely presented left orderable group $G=\langle\bar{c}\,|\,\mathcal{R}(\bar{c})\rangle$ with two elements $w_{1}(\bar{c}),w_{2}(\bar{c})\in G$ such that every homomorphic image of $G$ where $w_{1}(\bar{c})\not=w_{2}(\bar{c})$, has undecidable word problem. This group is torsion free. We may assume that values of all subwords appearing in $\mathcal{R}(\bar{c})\cup\{w_{1}(\bar{c}),w_{2}(\bar{c})\}$ are in $\bar{c}$. Applying Theorem 3.2 we see that $(\bar{c},\mathcal{R}(\bar{c})\cup\{w_{1}(\bar{c})\not=w_{2}(\bar{c})\})$ is a diagram without extensions witnessing WAP. Indeed, for every $(M,{\bf\bar{r}})\in[D(\bar{c})]$ the group $\langle\bar{c}\rangle$ (defined in $(M,{\bf\bar{r}})$) is a homomorphic image of $G$ and thus has undecidable word problem. $\Box$
Note that the proof of this corollary also demonstrates that the space $\mathcal{G}$ of all enumerable groups does not have generic groups. It also proves that the subspace of left orderable groups does not have generics too. These statements were major theorems of [6]. On the other hand we should emphasize that the idea of Darbinyan’s group used above is taken from [6]. We now give a new application. It is somehow opposite to the material of Section 2.3. Let $\mathbb{G}_{exp.n}$ denote the variety of groups of exponent $n$.
@@
Proof. Since the class of torsion groups of exponent $n$ is universally axiomatizable, this class forms a closed subspace of the logic space of enumerated groups. Furthermore, this axiomatization is computable. In particular every member of $\mathcal{B}_{T}$ is computably enumerable.
It is proved in [12] that there is a constant $C$ such that for every odd integer $n\geq C$, the following is true. Let a group $H$ from the variety $\mathbb{G}_{exp.n}$ be given by a countable set of generators $a_{1},a_{2},\ldots$ and a recursively enumerable set of defining relations $\mathcal{R}$ in these generators. Then $H$ is a subgroup of a 2-generated group $E$ finitely presented in $\mathbb{G}_{exp.n}$. Furthermore, it easily follows from the construction that the embedding $H\rightarrow E$ is computable. Indeed, it is shown in the proof of Corollary 1.2 of [12] that the group $H$ can be presented as a subgroup of $E=B(2,n)/P$ of the form $KP/P$, where
@@
Thus the map $a_{i}\rightarrow w_{i}$ defines the corresponding embedding in a computable way.
We now perform the Miller’s construction (i.e. from Proposition 5.2.13 of [6]) as follows. Fix recursively inseparable recursively enumerable subsets $M,N\subset\omega$. Let $A_{0}$ be the free abelian group on countably many generators $a_{i}$. Consider the following abelian quotient $A$ of $A_{0}$:
$A=\langle a_{1},a_{2},\ldots\,|\,a^{n}_{i}=1\,(i\in\omega)\,,\,a_{1}=a_{i}\mbox{ if }i\in M\mbox{ and }a_{2}=a_{j}\mbox{ if }j\in N\rangle.$
Let $\pi_{0}$ be the quotient map from $A_{0}$ to $A$. Applying the construction of Corollary 1.2 of [12] (described in the beginning of our proof) we computably embed $A$ into a 2-generated group $G$, which is finitely presented in $\mathbb{G}_{exp.n}$. Denote this embedding by $\phi$. Set $F=\{\phi(a_{1}),\phi(a_{2})\}\subset G$. Let us verify that, for any surjective homomorphism, say $\pi$ from $G$ to a group $H$ that is injective when restricted to $F$, we have that $H$ has undecidable word problem. Indeed, if the word problem of $H$ was decidable, then the set $L:=\{i\in\omega\,|\,(\pi\phi\pi_{0})(a_{i})=(\pi\phi\pi_{0})(a_{1})\}$ is computable, contains $M$, and is disjoint from $N$, contradicting the recursive inseparability of $M$ and $N$.

### v3

@@ 2.3 Nilpotent groups of prime odd exponent
Proof. We need some material concerning amalgamation bases in 2-step nilpotent groups, see Theorems 3.3 - 3.6 in [15]. It can be deduced from these theorems that a 2-step nilpotent group $G$ of exponent $p$ is an amalgamation base if and only if $G^{\prime}=Z(G)$. Below we only need sufficiency of this for finite groups, and the latter can be shown easier. Indeed, consider a finite $G\in\mathbb{G}_{2_{nil}p}$ with $G^{\prime}=Z(G)$. Note that any embedding of $G$ into any $H\in\mathbb{G}_{2_{nil}p}$ preserves every unary predicate $P$ with the property $H^{\prime}\leq P\leq Z(H)$ where $P$ on $G$ is defined by $P=G^{\prime}=Z(G)$. By Lemmas 2.2 and 2.3 of [1] the class of finite members of $\mathbb{G}_{2_{nil}p}$ in the language extended by $P$ with $H^{\prime}\leq P\leq Z(H)$, has the amalgamation property.
It now suffices to verify that every finite group $G_{0}\in\mathbb{G}_{2_{nil}p}$ extends to a finite group $G_{1}\in\mathbb{G}_{2_{nil}p}$ with $G^{\prime}_{1}=Z(G_{1})$. This can be again deduced from the amalgamation result of [1]. Indeed, consider $G_{0}$ with the predicate $P(G_{0})=Z(G_{0})$. Since $P(G_{0})$ is a vector space over $\mathbb{F}_{p}$ we decompose $P(G_{0})=G^{\prime}_{0}\oplus\langle e_{1},\ldots,e_{k}\rangle$ where $e_{1},\ldots,e_{k}$ is a basis of the complement of $G^{\prime}_{0}$. Now take a finite $H\in\mathbb{G}_{2_{nil}p}$ such that $H^{\prime}=Z(H)=\langle e_{1},\ldots,e_{k}\rangle$. The amalgamation process described in Lemmas 2.2 and 2.3 of [1] applied to $(G_{0},P)$ and $(H,P)$ gives the result. $\Box$
We do not know if the statement of this proposition holds for $c$-step nilpotent groups where $c>2$. The argument in the beginning of the proof together with [2] proves that if a finite group $G\in\mathbb{G}_{c_{nil}p}$ satisfies $\Gamma_{i}(G)=Z_{c+1-i}(G)$ for all $1<i\leq c$, then $G$ is an amalgamation base. Is it possible to embed every finite $H\in\mathbb{G}_{c_{nil}p}$ into some $G$ with this property?
Remark 2.12
The question formulated in the previous line has a strong connection with the following issue. Corollaries 1.3 - 1.5 of [2] state that the class of all finite $L^{+}$-structures from $\mathbb{G}^{P}_{c,p}$ has a Fraïssé limit, say $\mathsf{G}^{+}_{c,p}$, and it is $\aleph_{0}$-categorical, ultrahomogeneous and universal for countable groups from $\mathbb{G}_{c_{nil}p}$. It is natural to ask whether the reduct of this structure, say $\mathsf{G}_{c,p}$, to the group language is a generic structure from Theorem 2.9. The following sketch gives the affirmative under the assumption that the question formulated before the remark has the positive answer. According to Lemma 1.1 from [9] it suffices to prove that the reduct $\mathsf{G}_{c,p}$ belongs to every set of the form S, J, E from that lemma. The cases of J and E are rather easy by residual finiteness of finitely generated nilpotent groups and ultrahomogeneity of $\mathsf{G}^{+}_{c,p}$. We remind the reader that given a diagram $D$ the set ${\bf S}(D)$ consists of all enumerated groups $\mathsf{G}\in\mathcal{G}_{c_{nil}p}$ such that when $\mathsf{G}$ realizes $D$ then it also realizes some extension of $D$ witnessing WAP over $D$ in $\mathcal{B}_{c_{nil}p}$. To see this case it suffices to show that any finite subgroup $H_{1}<\mathsf{G}_{c,p}$ extends to a finite $H_{2}<\mathsf{G}_{c,p}$ such that every pair of finite extentions of $H_{2}$ is amalgamated over $H_{1}$. Take a finite $G\in\mathbb{G}_{c_{nil}p}$ that extends $H_{1}$ and satisfies $\Gamma_{i}(G)=Z_{c+1-i}(G)$ for all $1<i\leq c$. By ultrahomogeneity of $\mathsf{G}^{+}_{c,p}$, the unique expansion $G^{+}$ of $G$ to a structure from $\mathbb{G}^{P}_{c,p}$ has a copy in $\mathsf{G}^{+}_{c,p}$ over $H_{1}$. Let $H_{2}$ be the reduct of this copy to the group language. Since it is an amalgamation base, we can finish the argument. This, in particular, shows that the group $\mathsf{G}_{2,p}$ is a generic structure of the space $\mathcal{G}_{2_{nil}p}$ (see Proposition 2.11).
@@ 3 Strong undecidability implies non-WAP
In this section we give several group-theoretic properties such that $\mathcal{X}_{M,T}$ and the corresponding space of enumerated groups does not have generics. Thus, the section will focus on cases I - II of Section 1.2. On the other hand, the main technical tool, Theorem 3.2, works in much more general situations.
@@ 3.1 Completeness and Kuznetsov’s theorem
In Section 3 we assume that $M$ is an $\omega$-categorical structure in a language $L$ with a distinguished sort $\mathsf{Gp}$, and $T$ is an extension of $Th(M)$ in some language $L\cup\bar{r}$ satisfying the assumptions of Section 1.1 (i.e. our assumptions on $M$ are slightly stronger than in Section 2.1). When $(M,{\bf\bar{r}})$ is an expansion of $M$ in this language, we consider $\mathsf{Gp}$ as a structure with respect to those relations/functions of $L\cup\bar{r}$ which do not depend on elements of other sorts. Then any subset of $\mathsf{Gp}$ closed under operations of $\mathsf{Gp}$ is viewed as a substructure. If such a substructure is generated by some finite set $A$, then all valid atomic formulas depending on tuples from $A$ form a presentation of this substructure. Further notation corresponds to Section 2.1.
@@
Assume that $D(\bar{c})\in\mathcal{B}_{T}$ and an extension $D^{\prime}(\bar{c}^{\prime})\in\mathcal{B}_{T}$ is t-isolating for $\bar{c}$. Let $(M,{\bf\bar{r}})\in[D^{\prime}(\bar{c}^{\prime})]$. Assume that $Th(M,\bar{c}^{\prime})\cup D^{\prime}(\bar{c}^{\prime})\cup T$ is a computably enumerable set. Then in $(M,{\bf\bar{r}})$, the elements of $\bar{c}$ that belong to the sort $\mathsf{Gp}$ generate a recursively presented structure with decidable theory $\Xi_{\,\bar{c}\cap\mathsf{Gp},\bf\bar{r}}$.
Proof. We may assume that $\bar{c}$ contains all elements of the sort $\mathsf{Gp}$ which are distinguished in $(M,{\bf\bar{r}})$ as constants. Let us extend $D(\bar{c})$ by all atomic formulas with parameters just from $\bar{c}\cap\mathsf{Gp}$ that can be logically deduced from $Th(M,\bar{c}^{\prime})\cup D^{\prime}(\bar{c}^{\prime})\cup T$. We also include negations of such formulas when they are deduced from $Th(M,\bar{c}^{\prime})\cup D^{\prime}(\bar{c}^{\prime})\cup T$. The obtained set, say $\Sigma$, is computably enumerable. The set of atomic formulas from $\Sigma$ forms a computably enumerable presentation in the generators $\bar{c}\cap\mathsf{Gp}$ of the corresponding substructure generated in $\mathsf{Gp}$. Remember, that by t-isolation of $D^{\prime}(\bar{c}^{\prime})$ for $\bar{c}$, this presentation does not depend on the expansion $(M,{\bf\bar{r}})$.
The deciding algorithm is as follows. Let $\phi$ be an atomic formula with parameters from $\bar{c}\cap\mathsf{Gp}$. We verify $\phi\in\Sigma$ using the enumeration of $\Sigma$ obtained above. The algorithm stops when this procedure gives $\phi$ or $\neg\phi$ in $\Sigma$. If the first case does not hold, then in any member of $D^{\prime}(\bar{c}^{\prime})]$ the substructure generated by $\bar{c}\cap\mathsf{Gp}$ satisfies $\neg\phi$. If $\neg\phi$ does not belong to $\Sigma$, then $Th(M,\bar{c}^{\prime})\cup D^{\prime}(\bar{c}^{\prime})\cup T\cup\{\neg\phi\}$ is consistent. Any model of this theory does not have a reduct that is isomorphic to $(M,\bar{c}^{\prime})$ (since otherwise it defines a member of $D^{\prime}(\bar{c}^{\prime})]$). This contradicts $\omega$-categoricity of $Th(M,\bar{c}^{\prime})$. We see that the second case holds. $\Box$
The following theorem will be the main tool for proving the absence of generics in $\mathcal{G}_{T}$.
@@
Assume that $D(\bar{c})\in\mathcal{B}_{T}$ (where $\bar{c}$ includes all distinguished elements of $M$) satisfies the property that every extension $D^{\prime}(\bar{c}^{\prime})$ defines with $T\cup Th(M,\bar{c}^{\prime})$ a computably enumerable theory, and for every $(M,{\bf\bar{r}})\in[D(\bar{c})]$ the substructure of $\mathsf{Gp}$ generated by $\bar{c}\cap\mathsf{Gp}$ in $(M,{\bf\bar{r}})$ has undecidable theory $\Xi_{\,\bar{c}\cap\mathsf{Gp},\bf\bar{r}}$.
@@
Proof. Let $D^{\prime}(\bar{c}^{\prime})$ be an extension of $D(\bar{c})$. Let us show that it does not witness WAP. By Proposition 3.1, $D^{\prime}(\bar{c}^{\prime})$ is not t-isolating for $\bar{c}$. This means that there are $(M,{\bf\bar{r}})$, $(M,{\bf\bar{r}^{\prime}})\in[D^{\prime}(\bar{c}^{\prime})]$ and an atomic formula $\phi(\bar{c}_{0})$ with $\bar{c}_{0}\subseteq\bar{c}\cap\mathsf{Gp}$ such that $(M,{\bf\bar{r}})\models\phi(\bar{c}_{0})$ but $(M,{\bf\bar{r}^{\prime}})\models\neg\phi(\bar{c}_{0})$. We see that the two extensions of $D^{\prime}(\bar{c}^{\prime})$ by $\phi(\bar{c}_{0})$ and by $\neg\phi(\bar{c}_{0})$ (where domains of these extensions are enriched by values of all terms appeared in $\phi(\bar{c}_{0})$) cannot be amalgamated over $D(\bar{c})$. Therefore, $D^{\prime}(\bar{c}^{\prime})$ does not witness WAP. $\Box$
In the following corollary we assume Scenarios I - II of Section 1.2.
Corollary 3.3
Assume that $D(\bar{c})\in\mathcal{B}_{T}$ (where $\bar{c}$ includes all distinguished elements of $M$) satisfies the property that every extension $D^{\prime}(\bar{c}^{\prime})$ defines with $T\cup Th(M,\bar{c}^{\prime})$ a computably enumerable theory, and for every $(M,{\bf\bar{r}})\in[D(\bar{c})]$ the subgroup generated by $\bar{c}\cap\mathsf{Gp}$ in $(M,{\bf\bar{r}})$ has undecidable word problem.
Remark 3.4
Corollary 3.3 can be viewed as a counterpart of Theorem 1.1.6 of [8]. Indeed, by Theorem 1.2 from [9] both statements have the same conclusion: the absence of generic groups in $\mathcal{G}_{T}$. Their assumptions are also similar: in each of them undecidability of the word problem is involved in some hereditary way. As we will see below, both statements can be applied to the same objects with the same result. We also mention that Theorem 1.1.6 of [8] is a result of substantial effort: it appears in the final section of this long paper and is based on Theorem 1.1.3, another main theorem of the paper (among four main theorems which are stated by the authors in introduction).
We now give several applications of Corollary 3.3 for concrete spaces of enumerated groups. In these applications we are under Scenario I of Section 1.2. In particular, the structure $M$ is just $(\omega,1)$.
@@
In Proposition 5.2.13 of [8] the authors present a construction (of A. Darbinyan) of a finitely presented left orderable group $G=\langle\bar{c}\,|\,\mathcal{R}(\bar{c})\rangle$ with two elements $w_{1}(\bar{c}),w_{2}(\bar{c})\in G$ such that every homomorphic image of $G$ where $w_{1}(\bar{c})\not=w_{2}(\bar{c})$, has undecidable word problem. This group is torsion free. We may assume that values of all subwords appearing in $\mathcal{R}(\bar{c})\cup\{w_{1}(\bar{c}),w_{2}(\bar{c})\}$ are in $\bar{c}$. Applying Corollary 3.3 we see that $(\bar{c},\mathcal{R}(\bar{c})\cup\{w_{1}(\bar{c})\not=w_{2}(\bar{c})\})$ is a diagram without extensions witnessing WAP. Indeed, for every $(M,{\bf\bar{r}})\in[D(\bar{c})]$ the group $\langle\bar{c}\rangle$ (defined in $(M,{\bf\bar{r}})$) is a homomorphic image of $G$ and therefore has an undecidable word problem. Applying Theorem 1.2 from [9] we get the result. $\Box$
Note that the proof of this corollary also demonstrates that the space $\mathcal{G}$ of all enumerable groups does not have generic groups. It also proves that the subspace of left orderable groups does not have generics too. These statements are among the major theorems of [8]. On the other hand, we should emphasize that the idea of Darbinyan’s group used above is taken from [8]. We now give a new application. It is in some way opposite to the material in Section 2.3. Let $\mathbb{G}_{exp.n}$ denote the variety of groups of exponent $n$.
@@
Proof. Since the class of torsion groups of exponent $n$ is universally axiomatizable, this class forms a closed subspace of the logic space of enumerated groups. Furthermore, this axiomatization is computable. In particular, every member of $\mathcal{B}_{T}$ is computably enumerable.
It is proved in [14] that there is a constant $C$ such that for every odd integer $n\geq C$, the following is true. Let a group $H$ from the variety $\mathbb{G}_{exp.n}$ be given by a countable set of generators $a_{1},a_{2},\ldots$ and a recursively enumerable set of defining relations $\mathcal{R}$ in these generators. Then $H$ is a subgroup of a 2-generated group $E$ finitely presented in $\mathbb{G}_{exp.n}$. Furthermore, it easily follows from the description given in [14] that the embedding $H\rightarrow E$ is computable. Indeed, it is shown in the proof of Corollary 1.2 of [14] that the group $H$ can be presented as a subgroup of $E=B(2,n)/P$ of the form $KP/P$, where
@@
Thus, the map $a_{i}\rightarrow w_{i}$ defines the corresponding embedding in a computable way.
We now perform the Miller’s construction (i.e. from Proposition 5.2.13 of [8]) as follows. Fix recursively inseparable recursively enumerable subsets $N,N^{\prime}\subset\omega$ (i.e. there is no recursive set $W$ such that $N\subset W$ and $N^{\prime}\cap W=\emptyset$ ). Let $A_{0}$ be the free abelian group on countably many generators $a_{i}$. Consider the following abelian quotient $A$ of $A_{0}$:
$A=\langle a_{1},a_{2},\ldots\,|\,a^{n}_{i}=1\,(i\in\omega)\,,\,a_{1}=a_{i}\mbox{ if }i\in N\mbox{ and }a_{2}=a_{j}\mbox{ if }j\in N^{\prime}\rangle.$
Let $\pi_{0}$ be the quotient map from $A_{0}$ to $A$. Applying the construction of Corollary 1.2 of [14] (described in the beginning of our proof) we computably embed $A$ into a 2-generated group $G$, which is finitely presented in $\mathbb{G}_{exp.n}$. Denote this embedding by $\phi$. Set $F=\{\phi(a_{1}),\phi(a_{2})\}\subset G$. Let us verify that, for any surjective homomorphism, say $\pi$ from $G$ to a group $H$ that is injective when restricted to $F$, we have that $H$ has undecidable word problem. Indeed, if the word problem of $H$ was decidable, then the set $W:=\{i\in\omega\,|\,(\pi\phi\pi_{0})(a_{i})=(\pi\phi\pi_{0})(a_{1})\}$ is computable, contains $N$, and is disjoint from $N^{\prime}$, contradicting the recursive inseparability of $N$ and $N^{\prime}$.